OpenStudy (anonymous):

Simplify the given expression. square root negative 6 open parentheses 2 plus square root of negative 8

OpenStudy (anonymous):

didn't understand question can you write in mathematical form?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\sqrt{-6}\left( 2 + \sqrt{-8} \right)\]

OpenStudy (anonymous):

dont you want answer in points?

OpenStudy (anonymous):

6i(2+8i) =12i+48i^2 =12i-48 if thats what you want

OpenStudy (anonymous):

what do you mean? and no thats not one of my options

OpenStudy (anonymous):

sqrt(-6) can also be written as 6i..where i=sqrt(-1)

OpenStudy (anonymous):

12i-48i^2 isn't that your ans?

OpenStudy (anonymous):

\[2\sqrt{6} + 4\sqrt{3}\] \[-2\sqrt{6} + 4\sqrt{3}\] \[-4\sqrt{3} + 2i \sqrt{6}\] \[4\sqrt{3} + 2i \sqrt{6}\] these are my options

OpenStudy (candyme):

\[2\sqrt{-6} + \sqrt{-6}\sqrt{-8}\]\[2i \sqrt{6}+\sqrt{48}\]\[2i \sqrt6+4\sqrt{3}\]

OpenStudy (anonymous):

@Sarkar you have them under square root yet... \[\Large \sqrt{-6}(2+\sqrt{-8})=i\cdot \sqrt6(2+i\cdot \sqrt8)=\] \[\Large 2i\sqrt6 +i^2\sqrt{8\cdot 6}=2i\sqrt6+(-1)\cdot \sqrt{48}\] \[\Large 2i\sqrt6 -\sqrt{3\cdot 4^2}\] \[\Large 2i\sqrt6 -4\sqrt{3}\]

OpenStudy (anonymous):

but its not one of my options..

OpenStudy (anonymous):

oh yes my mistake..option C then

OpenStudy (anonymous):

ans is the third one

OpenStudy (anonymous):

then I made a mistake somewhere O_O

OpenStudy (anonymous):

hold on a second ... I'll grab calculator LOL

OpenStudy (candyme):

Oh, I'm sorry, in my answer I accidentally lost a negative

OpenStudy (anonymous):

no your ans is right

OpenStudy (anonymous):

2iunder root6−4under root3 can also be wriiten as −4underroot3+2iunderroot6

OpenStudy (anonymous):

didn't get?

OpenStudy (anonymous):

what do you call that ? O_O

OpenStudy (anonymous):

think its right@kreshnik

OpenStudy (anonymous):

yor 1st attenpt was right buddy.

OpenStudy (anonymous):

hmmm let me check

OpenStudy (anonymous):

... that's what I thought too @Sarkar .... @deadman340 I thought alike... but It's not in her options ! ;(

OpenStudy (anonymous):

3rd one is right as you did it earliere.

OpenStudy (anonymous):

@deadman340 . you're right, I'll erase my second one... thanks for assistance ;) Have a nice day ;)

OpenStudy (anonymous):

k dear